OpenStudy (anonymous):

Austin and Kaitlyn Kojan invested $195,000 in a business venture. If kaitlyn invested 2 1/4 times as much as Austin invested, how much money did Kaitlyn invest?

OpenStudy (kira_yamato):

2 1/4 = 9/4 Total is 9/4 + 1 = 13/4 Kaitlyn invested = 9/13 * 195000 = 135 000